Given are three vectors. Vector A of magnitude 7.10 units is in a direction θ = 31.0° above the positive x-axis. Vector B of magnitude 5.00 units is directed in the direction of the positive y-axis. Vector C of magnitude 6.60 units is in a direction ϕ = 37.0° clockwise from the negative y-axis. Find the magnitude and direction of 2 A − B + 3C. Find m agnitude and direction in degrees counterclockwise
Answer #1

Åg71 10=31° x 37 A= f 421 - - LM31, = 6086743-6571 0-5 @ : 66 Süs 37C-1) + 6665 (37)64) = -3.9727 -5.2719 2-6-6 8+= 2A-8+):